An installation requires 35KW and operates at a power factor of 0.45 from a 400v 3 phase supply what is the current taken by installation at the normal power factor 0.45

current taken at power factor 0.75

current taken at power factor of 0.9


1
Expert's answer
2021-01-27T10:20:35-0500

the line to neutral (phase) voltage, "V_{LN}=400/\sqrt{3}=230.94v"

three phase power is 35Kw, single phase power "=35\/3=11.6667Kw=11666.7w"

"Current , I =power/(voltage \times powerfactor)=11666.7/(230.94\times0.45)=112.26A"

current taken at power factor of 0.75"=11666.7/(230.94\times 0.75)=67.36A"

current taken at power factor of 0.9"=11666.7/(230.94\times 0.9)=56.13A"